OpenStudy (anonymous):

I dont understand this question... For what values of x does the function y= 1/4x -1 have y' = -4/49

OpenStudy (dumbcow):

\[(\frac{1}{4x-1})' = -\frac{4}{(4x-1)^2}\] set equal to -4/49, solve for x

OpenStudy (anonymous):

Thanks very much